In an experiment, 295 mg of copper is deposited when a current of 500 mA passes for 30 minutes. Find the electrochemical equivalent of copper?
1
Expert's answer
2018-05-21T08:10:08-0400
The Faraday’s law of electrolysis states
m=kIt
So, the electrochemical equivalent
k=m/It=(295×〖10〗^(-6) kg)/(500×〖10〗^(-3) A×30×60 s)=328.8×〖10〗^(-9) kg/s